Derive an expression for Van 't Hoff factor for a solute undergoing Association.
1
Expert's answer
2019-09-13T05:09:54-0400

Let's assume A associates to form "A_2"

So,

"\underset{\underset{n(1-a)}{n}}{2A}\to\ \underset{\underset{\frac{an}{2}}{0}}{A_2}"

Now if there were initially n moles of A and 0 moles of "A_2" then let's assume that ‘a’ times n moles associate so now moles of A=n(1-a) and that of "A_2" would be "\frac{na}{2}"

So now total no. of moles present


="n(1-a)+\frac{na}{2}=n-an+\frac{an}{2}\\n-\frac{an}{2}=n(1-\frac{a}{2})"

Van't Hoff factor = "\frac{total \ no.\ of\ moles\ after\ association}{Initial\ moles}=\frac{n(1-\frac{a}{2})}{n\\}\\"


"\ \ \ \ \ \ \ \ =(1-\frac{a}{2})"


where, "a" is degree of association